OpenStudy (anonymous):

Solve the system of equations and choose the correct answer from the list of options. x + y = −3 y = 2x + 2 A. 5/3 , 4/3 B. -5/3 , -4/3 C. -3/5 , -3/4 D. 3/4 , 3/5

OpenStudy (anonymous):

@pooja195

OpenStudy (mysticfury):

a good way to do these is to plug in the numbers :D the x coordinate is always first

OpenStudy (anonymous):

I've already plugged em in, x + (2x + 2) = -3

OpenStudy (anonymous):

i think the first one is -5/3, -4/3

OpenStudy (anonymous):

I just don't see where the fractions come in.

OpenStudy (anonymous):

neither do i

OpenStudy (anonymous):

And Mooshka-Skye thank you for your input but I am looking for an explanation not an answer

OpenStudy (anonymous):

There is no point in me advancing in my class if I don't understand what I am learning

OpenStudy (anonymous):

ok, and your welcome but im new by thirty minutes so yeah...

jhonyy9 (jhonyy9):

x + y = −3 y = 2x + 2 so the second equation is y=2x+2 - than just substitute this inside first equation in place of y and will get the value of x - ok ?

OpenStudy (anonymous):

I've substituted already, as said before I'm just not sure how to solve it. x + (2x + 2) = -3

jhonyy9 (jhonyy9):

so there are x+2x+2=-3 - now assume the same terms

jhonyy9 (jhonyy9):

courage this is easy

OpenStudy (anonymous):

Could it be B?

jhonyy9 (jhonyy9):

step by step please

OpenStudy (anonymous):

Alright just a sec

OpenStudy (anonymous):

forget it when i plugged in those fractions and worked backwards it didn't give me 41

OpenStudy (anonymous):

So it can not be B

jhonyy9 (jhonyy9):

please write here your work i like seeing where you have made it wrong - ok. ?

jhonyy9 (jhonyy9):

x + y = −3 y = 2x + 2 x+2x+2=-3 3x=-5 x=?

jhonyy9 (jhonyy9):

sorry but i dont understand your above words - please step by step continue please what i have wrote above

OpenStudy (anonymous):

-5/3, -1.7

jhonyy9 (jhonyy9):

so -5/3 is right but -1/7 how you get ?

OpenStudy (anonymous):

i meant -1.7 as in -5/3

OpenStudy (anonymous):

its easier for me to convert fractions to decimals then work with fractions alone

jhonyy9 (jhonyy9):

than you get for x= -5/3 so what will get for y=?

jhonyy9 (jhonyy9):

try with fractions and will get sure the proof of these results value

OpenStudy (anonymous):

Fractions aren't something that I can easily process. If you haven't noticed I'm not the brightest bulb.

OpenStudy (anonymous):

but -5/3 being x means it can only be B